OpenStudy (anonymous):

Solve x^2 - 4x = 5 x = 1 and x = -5 x = 1 and x = 5 x = -1 and x = 5 x = -1 and x = -5

OpenStudy (anonymous):

x^2-4x-5=0 (x-5)(x+1)=0 x-5=0---->x=5 x+1=0----->x=-1